# Classification of totally real number fields via their zeta function, regulator, and log unit lattice

### v1

@@ Abstract
keywords:
Log-unit lattice, regulator, Dedekind zeta function, arithmetic equivalence.
1 Acknowledgments
The author of this work is grateful to David Karpuk for suggesting the idea of studying log unit lattices of number fields, to Fatemeh Jalalvand for computing the minimum of the lattices in Example 1 and to Guillermo Mantilla-Soler for fruitful discussions and for suggesting Example 1. The author is also grateful to Kristaps Balodis, Erik Holmes, Fatemeh Jalalvand, Enrique Nunez Lon-Wo, Renate Scheidler and Ha Tran for fruitful discussions related to this paper.
@@ 2 Introduction
Now, since the regulator is computed by the covolume of the log unit lattice and the signature of its number field (see Definition 5), it is natural to wonder if having the same regulator implies having isometric log unit lattices. It turns out that for all of the examples mentioned above, their log unit lattices are not isometric, furthermore, their log unit lattices are not even similar.
In this paper, we address the following questions which are inspired by the above phenomena. All of our answers are conditional on the Weak Schanuel Conjecture. See 4.1.
@@ 2.1.1 On number fields with the same regulator and the linear independence of the residues
Assume 4.1 and let $K_{1},\,K_{2},\,\dots,\,K_{n}$ be totally real number fields all distinct and different from $\mathbb{Q}$ such that $\zeta_{K_{i}}\neq\zeta_{K_{j}}$ for every $i\neq j$. Then, the sets $\mathfrak{reg}:=\{\operatorname{reg}(K_{i})\}_{i=1}^{n}$ and $\mathfrak{res}:=\{\mathrm{res}_{s=1}\zeta_{K_{i}}(s)\}_{i=1}^{n}$ are linearly independent subsets of $\mathbb{C}$ over $\overline{\mathbb{Q}}$.
@@ 2.1.2 On the log unit lattice
Let us now state our results concerning the log unit lattice. Since the covolume of the log unit lattice is the regulator multiplied by an algebraic number (see Definition 5), it follows that the isometry class of the log unit lattice determines the projective point $\operatorname{reg}(K)]\in\mathbb{P}_{\overline{\mathbb{Q}}}(\mathbb{C})$. As a direct consequence of A, we find that, for totally real number fields, the isometry class conveys even more information:
@@
Assume 4.1. Let $K_{1}$ and $K_{2}$ be two totally real number fields and let $\Lambda_{K_{1}}$ and $\Lambda_{K_{2}}$ be the log unit lattices of $K_{1}$ and $K_{2}$, respectively. If $\Lambda_{K_{1}}$ and $\Lambda_{K_{2}}$ are isometric, then $\operatorname{reg}(K_{1})=\operatorname{reg}(K_{2})$. In particular, $\zeta_{K_{1}}=\zeta_{K_{2}}$ and $h_{K_{1}}=h_{K_{2}}$, where $h_{K_{1}}$ and $h_{K_{2}}$ are the class numbers of the corresponding number fields.
We conjecture that two totally real number fields with isometric log unit lattices are isomorphic, but the proof of this seems untractable at the moment. Nevertheless, we show in Example 1 a pair of totally real number fields with the same regulator but non-isometric and even non-similar log unit lattices. Hence, the isometry class of the log unit lattice conveys more information than the regulator about the isomorphism class of its field. This example can be found in [mantilla2015weak] and was suggested to the author of this work by Mantilla-Soler.
@@ 4 The genericity of the log unit lattice
Let $\mathcal{H}_{\overline{\mathbb{Q}}}\subset\mathbb{A}_{\overline{\mathbb{Q}}}^{n}$ be the the algebraic variety given by the hyperplane in the  $n$-dimensional $\overline{\mathbb{Q}}$-affine space satisfying the equation $\sum x_{i}=0$. Assume 4.1. Let $u\in\mathcal{O}_{K}^{\times}$ be a weak Minkowski unit, then $\mathrm{Log}(u)\in\mathcal{H}_{\mathbb{C}}$ is $\overline{\mathbb{Q}}$-generic.
@@ 5 The linear independence at $s=1$ of distinct Dedekind zeta functions
This section is dedicated to the proof of Theorem 5.3, which states, assuming 4.1, the linear independence of the residues at $s=1$ of distinct Dedekind zeta functions. We start by citing a special case of a result from [DeyGorlachKaihnsa2020] that we will use in the sequel. Afterwards, we will introduce the classical results of Gassman on arithmetic equivalence, these results allow us to reduce the proof to a representation theory statement. With these Tools in hand, we introduce some notation and then proceed with the proof of Theorem 5.3.
@@ 5.3 Proof of the linear independence
Assume 4.1 and let $K_{1},\,K_{2},\,\dots,\,K_{n}$ be totally real number fields all distinct and different from $\mathbb{Q}$ such that $\zeta_{K_{i}}\neq\zeta_{K_{j}}$ for every $i\neq j$. Then, the sets
@@
Note that by Proposition 3.4 and Remark 1 there exist nonzero constants $a_{1},a_{2},\dots,a_{n}\in\overline{\mathbb{Q}}$ such that
@@
Assume 4.1 and let $K_{1}$ and $K_{2}$ be two totally real number fields. If $\operatorname{reg}(K_{1})$ and $\operatorname{reg}(K_{2})$ are linearly dependent over the field of algebraic numbers $\overline{\mathbb{Q}}$, then $\zeta_{K_{1}}=\zeta_{K_{2}}$. Furthermore, $\operatorname{reg}(K_{1})=\operatorname{reg}(K_{2})$ if and only if $\zeta_{K_{1}}=\zeta_{K_{2}}$ and the class numbers $h_{K_{1}}$ and $h_{K_{2}}$ of $K_{1}$ and $K_{2}$, respectively, coincide.
@@ 6 On the log unit lattice as an invariant for number fields
We now investigate the strength of the isometry class and the similarity class of the log unit lattice as a classifying invariant for totally real number fields. We now prove Theorem 6.2 conditionally on 4.1. This theorem states that two real Galois number fields with similar log unit lattices are isomorphic.
@@ 6.1 Remarks on the isometry class of the log unit lattice of a totally real number field
Assume 4.1. Let $K_{1}$ and $K_{2}$ be two totally real number fields and let $\Lambda_{K_{1}}$ and $\Lambda_{K_{2}}$ be the log unit lattices of $K_{1}$ and $K_{2}$, respectively. If $\Lambda_{K_{1}}$ and $\Lambda_{K_{2}}$ are isometric, then $\operatorname{reg}(K_{1})=\operatorname{reg}(K_{2})$. In particular, $\zeta_{K_{1}}=\zeta_{K_{2}}$ and $h_{K_{1}}=h_{K_{2}}$, where $h_{K_{1}}$ and $h_{K_{2}}$ are the class numbers of the corresponding number fields.
@@ 6.2 The similarity class of the log unit lattice of a real Galois number field
Let us write $\mathrm{Gr}_{1}$, $\mathrm{Gr}_{2}$ and $\mathrm{Gr}_{M}$ for the corresponding Gram matrices. Note that by Remark 2 the units
@@
are weak Minkowski units of $K_{1}$ and $K_{2}$, respectively. Furthermore by Remark 1 and Lemma 3.1, for any Gram matrices $b_{K_{1}}$ and $b_{K_{2}}$ of $\Lambda_{K_{1}}$ and $\Lambda_{K_{2}}$, respectively, there exist matrices $C_{1}$ and $C_{2}$ and rational numbers $r_{1},r_{2}\in\mathbb{Q}$ such that

### v2

# Classification of totally real number fields via their zeta function, regulator, and log-unit lattice

@@ 1 Introduction
Now, since the regulator is computed by the covolume of the log unit lattice and the signature of its number field (see Definition 3.5), it is natural to wonder if having the same regulator implies having isometric log unit lattices. It turns out that for all of the examples mentioned above, their log unit lattices are not isometric, furthermore, their log unit lattices are not even similar.
In this paper, we address the following questions which are inspired by the above phenomena. All of our answers are conditional on the Weak Schanuel Conjecture. See 4.2.
@@ 1.1.1 On number fields with the same regulator and the linear independence of the residues
Assume 4.2 and let $K_{1},\,K_{2},\,\dots,\,K_{n}$ be totally real number fields all distinct and different from $\mathbb{Q}$ such that $\zeta_{K_{i}}\neq\zeta_{K_{j}}$ for every $i\neq j$. Then, the sets $\mathfrak{reg}:=\{\operatorname{reg}(K_{i})\}_{i=1}^{n}$ and $\mathfrak{res}:=\{\mathrm{res}_{s=1}\zeta_{K_{i}}(s)\}_{i=1}^{n}$ are linearly independent subsets of $\mathbb{C}$ over $\overline{\mathbb{Q}}$.
@@ 1.1.2 On the log unit lattice
Let us now state our results concerning the log unit lattice. Since the covolume of the log unit lattice is the regulator multiplied by an algebraic number (see Definition 3.5), it follows that the isometry class of the log unit lattice determines the projective point $\operatorname{reg}(K)]\in\mathbb{P}_{\overline{\mathbb{Q}}}(\mathbb{C})$. As a direct consequence of A, we find that, for totally real number fields, the isometry class conveys even more information:
@@
Assume 4.2. Let $K_{1}$ and $K_{2}$ be two totally real number fields and let $\Lambda_{K_{1}}$ and $\Lambda_{K_{2}}$ be the log unit lattices of $K_{1}$ and $K_{2}$, respectively. If $\Lambda_{K_{1}}$ and $\Lambda_{K_{2}}$ are isometric, then $\operatorname{reg}(K_{1})=\operatorname{reg}(K_{2})$. In particular, $\zeta_{K_{1}}=\zeta_{K_{2}}$ and $h_{K_{1}}=h_{K_{2}}$, where $h_{K_{1}}$ and $h_{K_{2}}$ are the class numbers of the corresponding number fields.
We conjecture that two totally real number fields with isometric log unit lattices are isomorphic, but the proof of this seems untractable at the moment. Nevertheless, we show in Example 6.2 a pair of totally real number fields with the same regulator but non-isometric and even non-similar log unit lattices. Hence, the isometry class of the log unit lattice conveys more information than the regulator about the isomorphism class of its field. This example can be found in [mantilla2015weak] and was suggested to the author of this work by Mantilla-Soler.
@@ 1.3 Outline of the paper:
2 Acknowledgments
The author of this work is grateful to David Karpuk for suggesting the idea of studying log unit lattices of number fields, to Fatemeh Jalalvand for computing the minimum of the lattices in Example 6.2 and to Guillermo Mantilla-Soler for suggesting Example 6.2. The author is also grateful to Kristaps Balodis, Erik Holmes, Fatemeh Jalalvand, Enrique Nunez Lon-Wo, Renate Scheidler and Ha Tran for fruitful discussions related to this paper.
@@ 4 The genericity of the log unit lattice
Let $\mathcal{H}_{\overline{\mathbb{Q}}}\subset\mathbb{A}_{\overline{\mathbb{Q}}}^{n}$ be the the algebraic variety given by the hyperplane in the  $n$-dimensional $\overline{\mathbb{Q}}$-affine space satisfying the equation $\sum x_{i}=0$. Assume 4.2. Let $u\in\mathcal{O}_{K}^{\times}$ be a weak Minkowski unit, then $\mathrm{Log}(u)\in\mathcal{H}_{\mathbb{C}}$ is $\overline{\mathbb{Q}}$-generic.
@@ 5 The linear independence at $s=1$ of distinct Dedekind zeta functions
This section is dedicated to the proof of Theorem 5.3, which states, assuming 4.2, the linear independence of the residues at $s=1$ of distinct Dedekind zeta functions. We start by citing a special case of a result from [DeyGorlachKaihnsa2020] that we will use in the sequel. Afterwards, we will introduce the classical results of Gassman on arithmetic equivalence, these results allow us to reduce the proof to a representation theory statement. With these Tools in hand, we introduce some notation and then proceed with the proof of Theorem 5.3.
@@ 5.3 Proof of the linear independence
Assume 4.2 and let $K_{1},\,K_{2},\,\dots,\,K_{n}$ be totally real number fields all distinct and different from $\mathbb{Q}$ such that $\zeta_{K_{i}}\neq\zeta_{K_{j}}$ for every $i\neq j$. Then, the sets
@@
Note that by Proposition 3.10 and Remark 4.6 there exist nonzero constants $a_{1},a_{2},\dots,a_{n}\in\overline{\mathbb{Q}}$ such that
@@
Assume 4.2 and let $K_{1}$ and $K_{2}$ be two totally real number fields. If $\operatorname{reg}(K_{1})$ and $\operatorname{reg}(K_{2})$ are linearly dependent over the field of algebraic numbers $\overline{\mathbb{Q}}$, then $\zeta_{K_{1}}=\zeta_{K_{2}}$. Furthermore, $\operatorname{reg}(K_{1})=\operatorname{reg}(K_{2})$ if and only if $\zeta_{K_{1}}=\zeta_{K_{2}}$ and the class numbers $h_{K_{1}}$ and $h_{K_{2}}$ of $K_{1}$ and $K_{2}$, respectively, coincide.
@@ 6 On the log unit lattice as an invariant for number fields
We now investigate the strength of the isometry class and the similarity class of the log unit lattice as a classifying invariant for totally real number fields. We now prove Theorem 6.4 conditionally on 4.2. This theorem states that two real Galois number fields with similar log unit lattices are isomorphic.
@@ 6.1 Remarks on the isometry class of the log unit lattice of a totally real number field
Assume 4.2. Let $K_{1}$ and $K_{2}$ be two totally real number fields and let $\Lambda_{K_{1}}$ and $\Lambda_{K_{2}}$ be the log unit lattices of $K_{1}$ and $K_{2}$, respectively. If $\Lambda_{K_{1}}$ and $\Lambda_{K_{2}}$ are isometric, then $\operatorname{reg}(K_{1})=\operatorname{reg}(K_{2})$. In particular, $\zeta_{K_{1}}=\zeta_{K_{2}}$ and $h_{K_{1}}=h_{K_{2}}$, where $h_{K_{1}}$ and $h_{K_{2}}$ are the class numbers of the corresponding number fields.
@@ 6.2 The similarity class of the log unit lattice of a real Galois number field
Let us write $\mathrm{Gr}_{1}$, $\mathrm{Gr}_{2}$ and $\mathrm{Gr}_{M}$ for the corresponding Gram matrices. Note that by Remark 6.3 the units
@@
are weak Minkowski units of $K_{1}$ and $K_{2}$, respectively. Furthermore by Remark 4.6 and Lemma 3.6, for any Gram matrices $b_{K_{1}}$ and $b_{K_{2}}$ of $\Lambda_{K_{1}}$ and $\Lambda_{K_{2}}$, respectively, there exist matrices $C_{1}$ and $C_{2}$ and rational numbers $r_{1},r_{2}\in\mathbb{Q}$ such that